# Ultracold Interactions between Ions and Polar Molecules

### v1

@@ 4 Reactive collisions and formation of molecular ions
In order to obtain systems with extremely high chemical reactivity that can be controlled on the quantum level, in addition to a suitably high polarizability, we also require energetically allowed target channels. To investigate if this is the case for the combination of Ba+ and NaK, we calculate the potential energy surfaces for all combinations of neutral and molecular reactants. As shown in Fig. 5 and 6, we indeed find several open reactive channels. We expect the production of ionic dimers NaK+, NaBa+, KBa+ as well as the atomic ions K+, Na+, Ba+ and all neutral constituents. This indicates that the system is highly suitable for studies of quantum chemistry at ultralow collision energies. However, the predicted energy release of the exoergic reactions is $\sim k_{B}\times 10^{4}~{}\textrm{K}$. In neutral molecule systems reactions on this energy scale are predominantly signaled by loss. Despite this complication, the rapid development in the field has allowed to obtain much more information about the collisional processes using kinetic energy spectrometry based on photo-ionization in conjunction with velocity-map imaging [84]. This is an example of how combining methodology from complementary fields can provide valuable new insights.
@@
In the present case, a hybrid ion-molecule apparatus would be equipped with an integrated Paul trap, such that, in principle, the expected ionic reaction products may be recaptured provided that the trap is suitable for this task. Typical Paul traps, including those employed in hybrid experiments, reach well depths on the order of $\sim k_{B}\times 10^{4}~{}\textrm{K}$ [85] which is comparable to the kinetic energy of the expected products. Moreover, the stability region of Paul traps for $\textrm{Ba}^{+}$ can be engineered to support the charge-to-mass ratios of the expected product ions which are similar to $\textrm{Rb}^{+}$ and $\textrm{Rb}_{2}^{+}$, species that have been sympathetically cooled by co-trapped $\textrm{Ba}^{+}$ [21, 86]. Furthermore, barrierless ion-neutral reactions provide an elegant way to synthesize and study polyatomic molecular ions in hybrid traps [87]. In this regard, similar systems may be realized either in successive ion-molecule collision experiments or by irradiating the ensembles with an ultraviolet photo-ionization beam, in analogy to recent studies of collisions between heteronuclear dimers [84]. Furthermore, it was shown that cooling of ions in a neutral buffer gas is extremely efficient in the classical collision regime and at high energies, even if the ions are not laser-coolable [51, 24]. In our case, ultracold and dense ensembles of three different neutral species, that is Na, K, and NaK can be used for this purpose. Therefore, it is likely that the molecular ions produced in ion-molecule reactions may be recaptured allowing for further manipulation and analysis. This includes sympathetic cooling [88], high-resolution mass spectrometry [86], and (Single shot) Doppler cooling thermometry (SS)DCT [89]. Ultimately, the produced molecules may also be investigated in high-precision measurements based on adapted quantum logic spectroscopy (QLS) methods [90, 91].
@@
5 Ions and high phase-space density polar molecules
Recent experiments capable of reaching ultralow ion-atom collision energies have also enabled first studies of interactions between atomic ions and loosely bound heteronuclear molecules [94]. Taking this concept further to a configuration where ions are immersed into a bath of polar molecules may enable novel polaronic effects due to the important role that the alignment of the molecules is expected to play in the dynamical evolution of such mixtures [80].
In the following we discuss the prospects for such experiments, based on the known properties of collisions between ions and neutral atoms and taking into account the influence of the ion’s electric field on the expected effective potentials. So far, we have restricted our configuration to the interaction of an ion with a single molecule prepared in its rovibronic ground state at large distances. In a next step, we will include the impact of the particle density and the internal energy structure of the molecules to derive the expected changes of the interaction potential when approaching the immersed ion.
@@ 5 Ions and high phase-space density polar molecules
In stark contrast to atoms with zero permanent electric dipole moment, whose collisions with ions are dominated by charge-induced dipole interaction, molecules exposed to the strong field in proximity of the ion become strongly aligned along the electric field axis. Their energy shifts are then dominated by the linear Stark effect. In the radially symmetric electric field, this gives rise to additional contributions to the effective potential due to the repulsive dipole-dipole interaction between pairs of molecules labeled as $i,j$ described by the following Hamiltonian:
@@
where $\varphi$ is the angle between the dipole moments $\hat{d}_{i,j}$ and $\vec{R}_{ij}$ the vector connecting the centers of mass of the two molecules.
@@
At distances where $\mathcal{E}\sim\mathcal{E}^{\star}$, the effective dipole moment is a sizable fraction of $d_{0}$ and the dominant ion-molecule interaction transitions from the attractive (in the rovibronic ground state) charge-induced dipole potential $-C_{4}/R^{4}$ in the long range to $U_{CD}\approx-d_{0}\mathcal{E}(R)\propto-d_{0}/R^{2}$ in vicinity of the ion where the dipole moments are almost fully aligned with respect to $\vec{\mathcal{E}}$. For example, in rovibronic ground-state NaK molecules the effective dipole moment amounts to $0.67~{}d_{0}$ at $\mathcal{E}=5~{}\mathcal{E}^{\star}$ as shown in Fig.4. Hence, the effective potential will be determined by the competition between dipole-dipole repulsion and ion-molecule interaction, such that the collective dipole-dipole term which is proportional to $d_{0}^{2}/R_{ij}^{3}$ will dominate at short range. At the point of equilibrium $R_{min}$, we expect a local minimum of the effective potential $U_{eff}$ experienced by each molecule of an ensemble contained within the interaction volume $V^{\star}$ of the ion.
@@
To confirm this expectation, we calculate $U_{eff}(R)$ by taking into account the relevant rotationally symmetric ion-molecule and anisotropic dipole-dipole interactions between molecules. For simplicity, we neglect the short-range van der Waals interaction $U_{vdW=}-C_{6}/R_{ij}^{6}$, and consider only next-neighbor contributions in an ensemble of $N$ molecules homogeneously distributed on a shell with radius $R_{min}$ centered around the ion, providing a lower bound on the repulsive interaction. We further take into account the finite induced electric dipole moment $d(\mathcal{E}\left(R\right))$ of the molecules (as shown in Fig.4) and the finite projection of $\vec{d}_{i}$ onto $\vec{d}_{j}$. The intermolecular distance $R_{ij}$ depends both on $R$ and the total number of molecules as $R_{ij}=2R\sin(\sqrt{\pi/N})$. Then, the effective dipole moment is $d_{eff}(R)=d(R)(1-\sin(\sqrt{\pi/N})^{2})$. Under these conditions and for an initial density of $n\approx 3\times 10^{10}~{}\textrm{cm}^{-3}$ we obtain the effective radial potential for a single molecule depicted in Fig.7. We find that it indeed exhibits a pronounced minimum with a depth of $U_{0}/k_{B}=56~{}\textrm{mK}$. To confirm that this effect persists with respect to variations of the molecular density, we repeat the calculation as a function of $N$. The results shown in Fig. 7 indicate that the potential minimum is robust with respect to the number of molecules, with its location shifting towards larger $R$ for higher neutral densities while at the same time the depth is decreasing, as illustrated in Fig.8.
@@
Figure 7: (left panel) Effective radial potential $U_{eff}/k_{B}$ per particle between a 138Ba+ ion and a quantum gas of 23Na39K molecules in their rovibronic ground state $\left|\textit{X}^{1}\Sigma^{+},\textrm{v}=0,J=0\right\rangle$ as a function of the ion-molecule distance $R$. We assume a moderate number density of $\textit{n}\approx 3\times 10^{10}~{}\textrm{cm}^{-3}$. The balance between attractive ion-neutral (charge-dipole, $\propto R^{-2}$, red dashed-dotted line) and repulsive molecule-molecule (dipole-dipole, $\propto R^{-3}$, blue dashed line) interactions gives rise to a global minimum at $R_{min}$, suggesting the formation of a polaron-like bound state on a shell at $R_{min}\approx 4.4\times 10^{-8}$ m. (center panel) Schematic three-dimensional illustration of a mesoscopic ion-molecule bound state. (right panel) Radial potential $U_{eff}/k_{B}$ (colorbar in mK) as a function of the molecule number $N$ within the interaction volume $V^{\star}$. A global minimum (largest negative value of $U_{eff}/k_{B}$) is expected for a wide range of $N$.
The calculated radial self-bound trapping frequency in the range $N=300$ to $1000$ would be on the order of $\omega_{sb}^{rad}/(2\pi)\sim 10~{}\textrm{MHz}$, e.g. for $N=350$, $\omega_{sb}^{rad}/(2\pi)\approx 35~{}\textrm{MHz}$ indicating that the molecules would be very strongly confined along the radial direction. The corresponding characteristic dipolar lengths and energies range between $a_{dd}=1.1\times 10^{-6}~{}\textrm{m}\approx 21100~{}a_{0}$ and $E_{d}/k_{B}=40~{}\textrm{nK}$ for $N=300$ and $a_{dd}=4.1\times 10^{-7}~{}\textrm{m}\approx 7700~{}a_{0}$ and $E_{d}/k_{B}=819~{}\textrm{nK}$ for $N=500$. These findings based on simplified static considerations provide a first indication that ions immersed into dense quantum gases of polar molecules may lead to the formation of many-body self-bound states (SBS).
Figure 8: Properties of the self-bound state (SBS) as a function of the number of molecules $N$ in NaK (top to bottom): SBS radius $R_{min}$, radial potential depth $U_{0}$, secular frequency $\omega_{rad}$, and intermolecular distance $R_{ij}$.
@@
Another approach to studying quantum-many body phenomena would be to immerse an ion into an ensemble of molecules confined to two dimensions. The latter can be realized, e.g. by utilizing 1D optical standing waves as a periodic trapping potential. In such setups the trapped molecules arrange themselves in layers with a regular spacing typically on the order of the wavelength of the dipole trap $\lambda$, as illustrated in Fig.9. This ensures very high trapping frequencies along the axis perpendicular to the trap layers such that the motional degrees of freedom are effectively constrained to a quasi-2D geometry. For an ion positioned at the center of a lattice site, that is for an ion-lattice spacing of $d_{lat}=0$, we calculate the effective potential of a molecule within the ensemble. As in the case of the 3D gas, we assume that the van der Waals interaction is negligible compared to contributions of $\hat{H}_{DD}$ and $\hat{H}_{CD}$, which could be achieved by tuning the s-wave scattering length to zero. In this geometry, we obtain $R_{ij}=2R\sin(\pi/N)$ and $d_{eff}(R)=d(R)\cos(\pi/N)$. Under such conditions, we again find local minima in the effective potential for a wide range of initial molecular densities, e.g. for $N=40$, where we obtain a well depth of $U_{0}/k_{B}=465~{}\textrm{mK}$, and $R_{min}=4\times 10^{-8}~{}\textrm{m}$. This indicates the existence of self-bound states in two dimensions.
Many of the discussed properties are reminiscent of polaron-type systems [4, 5, 6] and mesoscopic molecular ions [7] which have been studied extensively in the context of ion impurities in an atomic gas, but to the best of our knowledge no experimental realizations have yet been reported. It was suggested that such many-body states may form spontaneously at typical densities of $n=10^{14}~{}\textrm{cm}^{-3}$, with the binding energy being dissipated by phonon emission or by ion-atom TBR. The drastic enhancement of TBR at such densities would make the observation of mesoscopic molecule-atom states very challenging, but a similar mechanism could be present in both scenarios discussed here, aided by the expected suppression of TBR. Due to the presence of strong dipole-dipole interactions between the molecules, another way to release the binding energy could be by exciting collective surface vibration modes.
@@
Along similar lines, our results suggest that ions embedded in dense polar quantum gases are promising for observing intriguing transitions between, e.g., superfluid, Wigner crystal [95, 96] or supersolid phases [97]. For instance, the predicted quantum phase transition to a crystalline phase with a triangular lattice structure in 2D bosonic dipolar gases requires a dimensionless density of $n_{2D}a_{dd}^{2}>290$ [96]. While this is out of reach in existing contemporary experiments using polar molecules in 1D optical lattices, e.g. fermionic KRb [81], NaK has been suggested as a suitable molecule due to its larger PDM [98] and in the proposed system, an ion-molecule SBS with $N=350$ yields $n_{2D}a_{dd}^{2}>1.1\times 10^{4}$.
6 Ion-mediated shielding of reactive and inelastic collisions
@@
As shown in Fig. 8, typical separations between NaK molecules are $R_{ij}\approx 10~{}\textrm{nm}$. This is comparable to the characteristic van der Waals length of NaK $\bar{a}\approx 10~{}\textrm{nm}$ [99] and corresponds to strongly increased local densities approaching $n\approx 10^{18}~{}\textrm{cm}^{-3}$. In atom-ion mixtures, such high densities are inaccessible so far due to the extremely short lifetimes resulting from enhanced TBR [23, 25, 26, 27] and resonantly enhanced multi- photon ionization (REMPI) [100]. Ensembles of neutral molecules are also known to exhibit two-body and three-body losses with rates close to the universal limit attributed to sticky short-range collisions [37, 38, 35, 36]. Since collisional shielding techniques have not been demonstrated for densities in this region, we expect that the molecular ensembles would not be stable without a substantial polarization and the quasi-planar topology of the system.
@@
In the present scenario, the increase of local density is caused by the interaction of the molecules with the electric field of the ion. Due to their increasing alignment in the electric field mutual dipole-dipole repulsion eventually dominates at short range which allows for the formation of stable SBS. In potentials with the extremely high radial harmonic frequencies the bound molecules would confined to a shell. Due to the topologic similarity of a spherical surface with a planar disc (the corresponding spaces are homeomorphic up to a single point on the sphere), this confinement to a surface in 3D can be mapped to an equivalent geometry of polar molecules confined in a 2D plane, e.g., in a single site of a 1D optical standing wave. Since the radial electric field of the ion is locally parallel to the normal of any surface element on the shell, the configuration can be described by a homogeneous external electric field perpendicular to the plane containing molecules with individual effective dipole moments aligned parallel to the field and to each other. This situation where head-to-tail collisions are suppressed has been extensively studied with polar molecules, where efficient collisional shielding has been observed in degenerate Fermi gases of KRb [81, 44]. With the large degree of dipole moment alignment predicted for the SBS, we therefore expect highly efficient shielding of two and three-body processes. This should result in a suppression of sticky collisions, three-body losses and REMPI. As all molecules on the shell would be separated from the ion, we also expect that ion-neutral TBR will be efficiently suppressed. Notably, this ion-mediated shielding mechanism persists for or even benefits from very high local densities which are a prerequisite for entering the deep quantum-dominated regime of interactions.
7 Applications of hybrid ion-molecule systems
7.1 Control of collisional processes in molecular gases

### v2

@@ IV Reactive collisions and formation of molecular ions
[…]
rot. state $\sigma_{L}^{\textrm{NaK}}~{}(10^{-24}~{}\text{m}^{2})$ $K_{L}^{\textrm{Ba-NaK}}/K_{L}^{\textrm{Ba-Rb}}$
$\left|0,0\right\rangle$ $1.55$ $60$
$\left|1,0\right\rangle$ 00 00
$\left|1,\pm 1\right\rangle$ $0.85$ $32.9$
$\left|2,\pm 2\right\rangle$ $0.587$ $22.7$
$\left|3,\pm 2\right\rangle$ $0.0303$ $1.17$
[…]
@@
In order to obtain systems with extremely high chemical reactivity that can be controlled on the quantum level, in addition to a suitably high polarizability, we also require energetically allowed target channels [104]. To investigate if this is the case for the combination of Ba+ and NaK, we calculate the potential energy surfaces for all combinations of neutral and molecular reactants. As shown in Fig. 5 and 6, we indeed find several open reactive channels. We expect the production of ionic dimers NaK+, NaBa+, KBa+ as well as the atomic ions K+, Na+, Ba+ and all neutral constituents. This indicates that the system is highly suitable for studies of quantum chemistry at low and ultralow collision energies. However, the predicted energy release of the exoergic reactions is $\sim k_{B}\times 10^{4}~{}\textrm{K}$. In neutral molecule systems reactions on this energy scale are predominantly signaled by loss. Despite this complication, the rapid development in the field has allowed to obtain much more information about the collisional processes using kinetic energy spectrometry based on photo-ionization in conjunction with velocity-map imaging [43].
In the present case, a hybrid ion-molecule apparatus would be equipped with an integrated Paul trap, such that, in principle, the expected ionic reaction products may be recaptured provided that the trap is suitable for this task. Typical Paul traps, including those employed in hybrid experiments, reach well depths on the order of $\sim k_{B}\times 10^{4}~{}\textrm{K}$ [105] which is comparable to the kinetic energy of the expected products. Moreover, the stability region of Paul traps for $\textrm{Ba}^{+}$ can be engineered to support the charge-to-mass ratios of the expected product ions which are similar to $\textrm{Rb}^{+}$ and $\textrm{Rb}_{2}^{+}$, species that have been sympathetically cooled by co-trapped $\textrm{Ba}^{+}$ [23, 106]. Exploiting these properties, barrierless ion-neutral reactions may provide an elegant way to synthesize and study polyatomic molecular ions in hybrid traps [107]. In this regard, similar systems may be realized either in successive ion-molecule collision experiments or by irradiating the ensembles with an ultraviolet photo-ionization beam, in analogy to recent studies of collisions between heteronuclear dimers [43]. Furthermore, it was shown that cooling of ions in a neutral buffer gas is extremely efficient in the classical collision regime and at high energies, even if the ions are not laser-coolable [74, 26]. In our case, ultracold and dense ensembles of three different neutral species, that is Na, K, and NaK can be used for this purpose. Therefore, it is likely that the molecular ions produced in ion-molecule reactions may be recaptured allowing for further manipulation and analysis. This includes sympathetic cooling [108], high-resolution mass spectrometry [106], and (Single shot) Doppler cooling thermometry (SS)DCT [109]. Ultimately, the produced molecules may also be investigated in high-precision measurements based on adapted quantum logic spectroscopy (QLS) methods [110, 111].
V Ions and high phase-space density polar molecules
V.1 Role of dipole-dipole interactions
Recent experiments capable of reaching ultralow ion-atom collision energies have also enabled first studies of interactions between atomic ions and loosely bound heteronuclear molecules [112]. Taking this concept further to a configuration where ions are immersed into a bath of polar molecules may allow exploring a new regime of interactions due to the important role that the alignment of the molecules is expected to play in the dynamical evolution of such mixtures [70].
In the following, we discuss the prospects for such experiments, based on the known properties of collisions between ions and neutral atoms and taking into account the influence of the ion’s electric field on the expected effective potentials. Here, we include the impact of the particle density and the internal energy structure of the molecules in their rovibronic ground state to derive the expected changes of the interaction potential when approaching the immersed ion.
@@ V.1 Role of dipole-dipole interactions
In stark contrast to atoms, whose collisions with ions are dominated by charge-induced dipole interaction $U_{CID}=-C_{4}/R^{4}$, molecules in proximity of the ion become strongly aligned along the electric field axis. Their energy shifts are then dominated by the linear Stark effect. In the radially symmetric electric field, this gives rise to additional contributions to the effective potential due to the repulsive dipole-dipole interaction between pairs of molecules labeled as $i,j$ described by the following Hamiltonian:
@@
where $\varphi$ is the angle between the dipole moments $\hat{d}_{i,j}$ and $\vec{R}_{ij}$ the vector connecting the centers of mass of the two molecules. At distances where $\mathcal{E}\sim\mathcal{E}^{\star}$, the effective dipole moment is a sizable fraction of $d_{0}$ and the dominant ion-molecule interaction transitions from the attractive (in the rovibronic ground state) $U_{CID}$ in the long range to $U_{CD}\approx-d_{0}\mathcal{E}(R)\propto-d_{0}/R^{2}$ in vicinity of the ion where the dipole moments are almost fully aligned with respect to $\vec{\mathcal{E}}$. For example, in rovibronic ground-state NaK molecules the effective dipole moment amounts to $0.67~{}d_{0}$ at $\mathcal{E}=5~{}\mathcal{E}^{\star}$ as shown in Fig.2.
@@
As a potential scenario where the competition between attractive and repulsive interactions would play an important role we consider an ensemble of homogeneously distributed molecules at a distance $R$ around the ion. In this case, while the molecules approach the ion, the collective dipole-dipole term which is proportional to $d_{0}^{2}/R_{ij}^{3}$ will dominate at short range. For the sake of simplicity, we have neglected the short-range van der Waals interaction $U_{vdW=}-C_{6}/R_{ij}^{6}$, and consider only next-neighbor contributions between the molecules. We further take into account the finite induced electric dipole moment $d(\mathcal{E}\left(R\right))$ of the molecules (as shown in Fig.2) and the finite projection of $\vec{d}_{i}$ onto $\vec{d}_{j}$. The intermolecular distance is $R_{ij}=2R\sin(\sqrt{\pi/N})$ and the effective dipole moment is $d_{eff}(R)=d(R)(1-\sin(\sqrt{\pi/N})^{2})$. Under these conditions and for an initial density of $n\approx 4\times 10^{10}~{}\textrm{cm}^{-3}$ we obtain the effective radial potential $U_{eff}$ for a single molecule depicted in Fig.7. We find that the competition of repulsive intermolecular and attractive ion-molecule interactions in the considered configuration gives rise to a local minimum around $R_{min}\approx 40~{}\textrm{nm}$ which is shifted towards larger $R$ for an increasing number of molecules while its depth $U_{eff}/k_{B}\approx 80~{}\textrm{mK}$ decreases for larger $N$.
@@
Figure 7: Effective radial potential $U_{eff}/k_{B}$ per particle between a 138Ba+ ion and a quantum gas of 23Na39K molecules in their rovibronic ground state $\left|\textit{X}^{1}\Sigma^{+},\textrm{v}=0,J=0\right\rangle$ as a function of the ion-molecule distance $R$. We assume a moderate number density of $\textit{n}\approx 4\times 10^{10}~{}\textrm{cm}^{-3}$. The balance between attractive ion-neutral (charge-dipole, $\propto R^{-2}$, red dashed-dotted line) and repulsive molecule-molecule (dipole-dipole, $\propto R^{-3}$, blue dashed line) interactions gives rise to a global minimum at $R_{min}$.
@@
For simplicity, we have focused on an electrostatic view and omitted dynamical aspects such as a mechanism to dissipate kinetic energy once $R_{min}$ has been reached or a detailed stability analysis, which would be important for future realistic theoretical or experimental studies of such shell-like configurations in a 3D bulk gas. A potentially more accessible approach to studying the interplay of several simultaneously active interactions and their impact on the dynamics of a many-body system would be to immerse an ion into an ensemble of molecules confined to two dimensions. The latter can be realized, e.g. by utilizing 1D optical standing waves as a periodic trapping potential. In such setups the trapped molecules arrange themselves in layers with a regular spacing typically on the order of the wavelength of the dipole trap $\lambda$. This ensures very high trapping frequencies along the axis perpendicular to the trap layers such that the motional degrees of freedom are effectively constrained to a quasi-2D geometry. For an ion positioned at the center of a lattice site, that is for an ion-lattice spacing of $d_{lat}=0$, we calculate the effective potential of a molecule within the ensemble. In this geometry, we obtain $R_{ij}=2R\sin(\pi/N)$ and $d_{eff}(R)=d(R)\cos(\pi/N)$. Under such conditions, we again find local minima in the effective potential for a wide range of initial molecular densities.
@@
Such effects are an interesting field for future investigations, providing connections to polaron-type systems [3, 4, 5, 7, 8] and mesoscopic molecular ions [6] which have been studied extensively in the context of ion impurities in an atomic gas. One possible direction would be to explore potential transitions between, e.g., superfluid, Wigner crystal [113, 114] or supersolid phases [115] induced by embedding ions in dense polar quantum gases. For instance, the predicted quantum phase transition to a crystalline phase with a triangular lattice structure in 2D bosonic dipolar gases requires a dimensionless density of $n_{2D}a_{dd}^{2}>290$ [114]. While this is out of reach in existing contemporary experiments using polar molecules in 1D optical lattices, e.g. fermionic KRb [98], NaK has been suggested as a suitable molecule due to its larger PDM [116]. Under the assumption that, e.g. $N=500$ molecules can be arranged in a stable configuration as the one discussed above in a realistic experiment, the system would yield a dimensionless density of $n_{2D}a_{dd}^{2}\approx 3\times 10^{4}$ which would make it promising for observing such a phase transition.
V.2 Ion-mediated shielding of reactive and inelastic collisions
@@
In the following, we consider the impact of high molecular densities expected for the configurations discussed in the previous Section assuming they may be realized. For the example above, typical separations between NaK molecules would be $R_{ij}\approx 10~{}\textrm{nm}$. This is comparable to the characteristic van der Waals length of NaK $\bar{a}\approx 10~{}\textrm{nm}$ [71] and corresponds to strongly increased local densities approaching $n\approx 10^{18}~{}\textrm{cm}^{-3}$. In atom-ion mixtures, such high densities are inaccessible so far due to the extremely short lifetimes resulting from enhanced TBR [25, 27, 28, 29] and resonantly enhanced multi-photon ionization (REMPI) [117]. Ensembles of neutral molecules are also known to exhibit two-body and three-body losses with rates close to the universal limit attributed to sticky short-range collisions [37, 44, 38, 39]. Since collisional shielding techniques have not been demonstrated for densities in this region, we expect that the molecular ensembles would not be stable without a substantial polarization and the quasi-planar topology of the system.
@@
In the present scenario, the increase of local density is caused by the interaction of the molecules with the electric field of the ion. Due to their increasing alignment in the electric field mutual dipole-dipole repulsion eventually dominates at short range, which would result in the local minimum of the effective potential shown in Fig. 7 corresponding to a tightly confining shell. Due to the topologic similarity of a spherical surface with a planar disc (the corresponding spaces are homeomorphic up to a single point on the sphere), this confinement to a surface in 3D can be mapped to an equivalent geometry of polar molecules confined in a 2D plane, e.g., in a single site of a 1D optical standing wave. Since the radial electric field of the ion is locally parallel to the normal of any surface element on the shell, the configuration can be described by a homogeneous external electric field perpendicular to the plane containing molecules with individual effective dipole moments aligned parallel to the field and to each other. This situation where head-to-tail collisions are suppressed has been extensively studied with polar molecules, where efficient collisional shielding has been observed in degenerate Fermi gases of KRb [98, 52]. With the large degree of dipole moment alignment present in the discussed configuration, we therefore expect highly efficient shielding of two and three-body processes. This should result in a suppression of sticky collisions, TBR losses and REMPI. As all molecules on the shell would be separated from the ion, we also expect that ion-neutral TBR would be efficiently suppressed. Notably, this ion-mediated shielding mechanism would persist for or even benefit from very high local densities which are a prerequisite for entering the deep quantum-dominated regime of interactions. Similarly, we also expect efficient ion-mediated shielding for the case of dense molecular gases in a quasi-2D geometry achieved by confinement in a 1D optical lattice as discussed above.
VI Applications of hybrid ion-molecule systems
VI.1 Control of collisional processes in molecular gases